# SLRQA: A Sparse Low-Rank Quaternion Model for Color Image Processing with Convergence Analysis

### v1

@@ 4 Convergence analysis
Suppose A2,A3 in Assumption 4.1 hold, the sequence $\{({\mathbf{\dot{X}}}_{k+1},\mathbf{\dot{W}}_{k+1},\mathbf{\dot{\Lambda}}_{k+1})\}_{k\geq 0}$ generated by (17) satisfies:
$\frac{1}{\beta\mu}\|\mathbf{\dot{\Lambda}}_{k+1}\|^{2}_{\mathrm{F}}\leq\theta_{0}\|\Delta\mathbf{\dot{W}}_{k}\|^{2}_{\mathrm{F}}+\theta_{1}\|\Delta\mathbf{\dot{W}}_{k+1}\|^{2}_{\mathrm{F}}+\theta_{2}\|\mathcal{B}^{\#}(\Delta\mathbf{\dot{\Lambda}}_{k})\|^{2}_{\mathrm{F}}-\theta_{2}\|\mathcal{B}^{\#}(\Delta\mathbf{\dot{\Lambda}}_{k+1})\|^{2}_{\mathrm{F}},$ (24)
$\mathcal{L}_{\beta}({\mathbf{\dot{X}}}_{k+1},\mathbf{\dot{W}}_{k+1},\mathbf{\dot{\Lambda}}_{k+1})+\|\Delta\mathbf{\dot{W}}_{k+1}\|_{B_{k}-r\theta_{1}\mathbf{I}}^{2}+\frac{L_{1,k}}{2}\|\Delta{\mathbf{\dot{X}}}_{k+1}\|^{2}_{\mathrm{F}}+\frac{r-1}{\beta\mu}\|\Delta\mathbf{\dot{\Lambda}}_{k+1}\|^{2}_{\mathrm{F}}$
$+r\theta_{2}\|\mathcal{B}^{\#}\Delta\mathbf{\dot{\Lambda}}_{k+1}\|^{2}_{\mathrm{F}}\leq\mathcal{L}_{\beta}({\mathbf{\dot{X}}}_{k},\mathbf{\dot{W}}_{k},\mathbf{\dot{\Lambda}}_{k})+r\theta_{0}\|\Delta\mathbf{\dot{W}}_{k}\|^{2}_{\mathrm{F}}+r\theta_{2}\|\mathcal{B}^{\#}(\Delta\mathbf{\dot{\Lambda}}_{k})\|^{2}_{\mathrm{F}},$ (25)
where $r>1,\theta_{0},\theta_{1}$ are defined in (21), and $\theta_{2}:=\frac{|1-\mu|}{\beta\mu\lambda_{+}^{\mathcal{B}^{\#}\mathcal{B}}\rho(\mu)}.$
@@
$\mathbf{\dot{G}}_{k+1}:=-L_{2,k}\Delta\mathbf{\dot{W}}_{k+1}-\nabla g(\mathbf{\dot{W}}_{k}).$ (26)
It follows that $\Delta\mathbf{\dot{G}}_{k+1}=L_{2,k-1}\Delta\mathbf{\dot{W}}_{k}-L_{2,k}\Delta\mathbf{\dot{W}}_{k+1}+\nabla g(\mathbf{\dot{W}}_{k-1})-\nabla g(\mathbf{\dot{W}}_{k}).$ By the triangle inequality, the following inequality holds:
$\|\Delta\mathbf{\dot{G}}_{k+1}\|_{\mathrm{F}}\leq\|\nabla g(\mathbf{\dot{W}}_{k-1})-\nabla g(\mathbf{\dot{W}}_{k})\|_{\mathrm{F}}+L_{2,k}\|\Delta\mathbf{\dot{W}}_{k+1}\|_{\mathrm{F}}+L_{2,k-1}\|\Delta\mathbf{\dot{W}}_{k}\|_{\mathrm{F}}.$
By A2 in Assumption 4.1, $\nabla g(\mathbf{\dot{W}})$ is $L_{g}$ Lipschitz continuous and $q_{2}=\sup_{k\geq 0}L_{2,k}<\infty,$ we have $\|\Delta\mathbf{\dot{G}}_{k+1}\|_{\mathrm{F}}\leq(L_{g}+q_{2})\|\Delta\mathbf{\dot{W}}_{k}\|_{\mathrm{F}}+q_{2}\|\Delta\mathbf{\dot{W}}_{k+1}\|_{\mathrm{F}}.$ Hence it follows that:
@@
Let $\mathcal{R}({\mathbf{\dot{X}}},\mathbf{\dot{W}},\mathbf{\dot{\Lambda}},\mathbf{\dot{W}}^{\prime},\mathbf{\dot{\Lambda}}^{\prime}):=\mathcal{L}_{\beta}({\mathbf{\dot{X}}},\mathbf{\dot{W}},\mathbf{\dot{\Lambda}})+r\theta_{0}\|\mathbf{\dot{W}}-\mathbf{\dot{W}}^{\prime}\|^{2}_{\mathrm{F}}+r\theta_{2}\|\mathcal{B}^{\#}(\mathbf{\dot{\Lambda}}-\mathbf{\dot{\Lambda}}^{\prime})\|^{2}_{\mathrm{F}}$. The merit function $\mathcal{R}_{k}$ is defined by
$\mathcal{R}_{k}:=\mathcal{R}({\mathbf{\dot{X}}}_{k},\mathbf{\dot{W}}_{k},\mathbf{\dot{\Lambda}}_{k},\mathbf{\dot{W}}_{k-1},\mathbf{\dot{\Lambda}}_{k-1})=\mathcal{L}_{\beta}({\mathbf{\dot{X}}}_{k},\mathbf{\dot{W}}_{k},\mathbf{\dot{\Lambda}}_{k})+r\theta_{0}\|\Delta\mathbf{\dot{W}}_{k}\|^{2}_{\mathrm{F}}+r\theta_{2}\|\mathcal{B}^{\#}\Delta\mathbf{\dot{\Lambda}}_{k}\|^{2}_{\mathrm{F}}.$ (30)
@@
$\leq\mathcal{R}_{k+1}+\|\Delta\mathbf{\dot{W}}_{k+1}\|_{B_{k}-r(\theta_{1}+\theta_{0})\mathbf{I}}^{2}+\frac{L_{1,k}}{2}\|\Delta{\mathbf{\dot{X}}}_{k+1}\|^{2}_{\mathrm{F}}+\frac{r-1}{\beta\mu}\|\Delta\mathbf{\dot{\Lambda}}_{k+1}\|^{2}_{\mathrm{F}}\leq\mathcal{R}_{k}\leq\mathcal{R}_{k_{0}},$
@@
$-\frac{1}{2\beta}\left\|\mathbf{\dot{\Lambda}}_{k+1}\right\|_{\mathrm{F}}^{2}+(r\theta_{0}+a)\|\Delta\mathbf{\dot{W}}_{k+1}\|_{\mathrm{F}}^{2}+a(\|\Delta{\mathbf{\dot{X}}}_{k+1}\|_{\mathrm{F}}^{2}+\|\Delta\mathbf{\dot{\Lambda}}_{k+1}\|_{\mathrm{F}}^{2})+r\theta_{2}\|\mathcal{B}^{\#}\Delta\mathbf{\dot{\Lambda}}_{k+1}\|_{\mathrm{F}}^{2}\leq\mathcal{R}_{k_{0}}.$
@@
By the convexity of $\|\cdot\|^{2}$, we have $\lambda_{+}^{\mathcal{B}^{\#}\mathcal{B}}\mu^{2}\|\mathbf{\dot{\Lambda}}_{k+1}\|_{\mathrm{F}}^{2}\leq\frac{\mu^{2}}{\rho(\mu)}\|\mathbf{\dot{G}}_{k+1}\|_{\mathrm{F}}^{2}+|1-\mu|\|\mathcal{B}^{\#}\Delta\mathbf{\dot{\Lambda}}_{k+1}\|_{\mathrm{F}}^{2}.$ According to definition (26), we have $\|\mathbf{\dot{G}}_{k+1}\|_{\mathrm{F}}^{2}\leq 2q_{2}^{2}\|\Delta\mathbf{\dot{W}}_{k+1}\|_{\mathrm{F}}^{2}+2\|\nabla g(\mathbf{\dot{W}}_{k})\|_{\mathrm{F}}^{2}\leq 2(q_{2}+L_{g})^{2}\|\Delta\mathbf{\dot{W}}_{k+1}\|_{\mathrm{F}}^{2}+2\|\nabla g(\mathbf{\dot{W}}_{k+1})\|_{\mathrm{F}}^{2}.$ It follows that:
$-\frac{1}{2\beta}\|\mathbf{\dot{\Lambda}}_{k+1}\|_{\mathrm{F}}^{2}\geq-\theta_{3}\|\nabla g(\mathbf{\dot{W}}_{k+1})\|_{\mathrm{F}}^{2}-\theta_{4}\|\Delta\mathbf{\dot{W}}_{k+1}\|_{\mathrm{F}}^{2}-\theta_{5}\|\mathcal{B}^{\#}\Delta\mathbf{\dot{\Lambda}}_{k+1}\|_{\mathrm{F}}^{2},$ (34)
@@
$\theta_{3}:=\frac{1}{\beta\rho(\mu)\lambda_{+}^{\mathcal{B}^{\#}\mathcal{B}}},\quad\theta_{4}:=\frac{(q_{2}+L_{g})^{2}}{\beta\rho(\mu)\lambda_{+}^{\mathcal{B}^{\#}\mathcal{B}}},\quad\theta_{5}:=\frac{|1-\mu|}{2\beta\mu^{2}\lambda_{+}^{\mathcal{B}^{\#}\mathcal{B}}}.$
@@
$f({\mathbf{\dot{X}}}_{k+1})+\frac{1}{2\tau^{2}}\|\mathcal{P}_{\Omega}(\mathcal{A}({\mathbf{\dot{X}}}_{k}))-\mathcal{P}_{\Omega}(\mathbf{\dot{Y}})\|_{\mathrm{F}}^{2}+\frac{\beta}{2}\left\|\mathcal{C}({\mathbf{\dot{X}}}_{k+1})+\mathcal{B}(\mathbf{\dot{W}}_{k+1})-\mathbf{\dot{B}}+\frac{\mathbf{\dot{\Lambda}}_{k+1}}{\beta}\right\|_{\mathrm{F}}^{2}$ (35)
$+a(\|\Delta{\mathbf{\dot{X}}}_{k+1}\|_{\mathrm{F}}^{2}+\|\Delta\mathbf{\dot{\Lambda}}_{k+1}\|_{\mathrm{F}}^{2})+(r\theta_{0}+a-\theta_{4})\|\Delta\mathbf{\dot{W}}_{k+1}\|_{\mathrm{F}}^{2}+(r\theta_{2}-\theta_{5})\|\mathcal{B}^{\#}\Delta\mathbf{\dot{\Lambda}}_{k+1}\|_{\mathrm{F}}^{2}$
$\leq\mathcal{R}_{k_{0}}-\inf_{\mathbf{\dot{W}}}\left\{g(\mathbf{\dot{W}})-\theta_{3}\|\nabla g(\mathbf{\dot{W}})\|_{\mathrm{F}}^{2}\right\}.$
According to (19) in Assumption A2, setting $\mathbf{\dot{W}}_{1}=\mathbf{\dot{W}}-\delta\nabla g(\mathbf{\dot{W}})$ and $\mathbf{\dot{W}}_{2}=\mathbf{\dot{W}}$, it follows that $g(\mathbf{\dot{W}}_{k}-\delta\nabla g(\mathbf{\dot{W}}_{k}))\leq g(\mathbf{\dot{W}}_{k})-(\delta-\frac{L_{g}\delta^{2}}{2})\|\nabla g(\mathbf{\dot{W}}_{k})\|_{\mathrm{F}}^{2}.$ Since $g$ is bounded from below, there exist $M$ such that
$-M<\inf\{g(\mathbf{\dot{W}})-(\delta-\frac{L_{g}\delta^{2}}{2})\|\nabla g(\mathbf{\dot{W}})\|_{\mathrm{F}}^{2}:\mathbf{\dot{W}}\in\mathbb{H}^{m\times n}\}.$ (36)
We choose $\delta=\frac{1}{L_{g}}$. According to Assumption A5, we have $\theta_{3}<\frac{1}{2L_{g}}=\delta-\frac{L_{g}\delta^{2}}{2}.$ Since $r>1$ and $\mu\in(0,2),$ according to the denifition of $\theta_{0},\theta_{2},\theta_{4}$ and $\theta_{5}$, it holds that $r\theta_{2}-\theta_{5}>0$ and $r\theta_{0}+a-\theta_{4}>0$. It follows from (35) that
$f({\mathbf{\dot{X}}}_{k+1})+\frac{1}{2\tau^{2}}\|\mathcal{P}_{\Omega}(\mathcal{A}({\mathbf{\dot{X}}}_{k}))-\mathcal{P}_{\Omega}(\mathbf{\dot{Y}})\|_{\mathrm{F}}^{2}$ (37)
@@
Since $f$ is coercive, the sequence $\{{\mathbf{\dot{X}}}_{k}\}_{k\geq k_{0}}$ is bounded and hence $\{\mathcal{C}{\mathbf{\dot{X}}}_{k}\}_{k\geq k_{0}}$ is bounded. According to (37), $\Delta\mathbf{\dot{\Lambda}}_{k+1}$ is bounded. By the $\mathbf{\dot{\Lambda}}$ update in (17), we have $\mathcal{B}\mathbf{\dot{W}}_{k+1}=\frac{1}{\beta\mu}\Delta\mathbf{\dot{\Lambda}}_{k+1}-\mathcal{C}{\mathbf{\dot{X}}}_{k+1}+\mathbf{\dot{B}}.$ Since $\{\mathcal{C}{\mathbf{\dot{X}}}_{k}\}_{k\geq k_{0}}$ is bounded and $\mathcal{B}^{\#}\mathcal{B}$ is full rank, it follows that $\{\mathbf{\dot{W}}_{k}\}_{k\geq k_{0}}$ is bounded. From the fact that $\frac{\beta}{2}\left\|\mathcal{C}({\mathbf{\dot{X}}}_{k+1})+\mathcal{B}(\mathbf{\dot{W}}_{k+1})-\mathbf{\dot{B}}+\frac{\mathbf{\dot{\Lambda}}_{k+1}}{\beta}\right\|_{\mathrm{F}}^{2}$ is bounded, it follows that $\{\mathbf{\dot{\Lambda}}_{k}\}_{k\geq k_{0}}$ is bounded. As a consequence, $\{{\mathbf{\dot{X}}}_{k}\}_{k\geq 1},\{\mathbf{\dot{W}}_{k}\}_{k\geq 1},\{\mathbf{\dot{\Lambda}}_{k}\}_{k\geq 1}$ is bounded. ∎
@@
$\mathbf{\dot{D}}_{{\mathbf{\dot{X}}}_{k+1}}$ $:=\mathcal{C}^{\#}\Delta\mathbf{\dot{\Lambda}}_{k+1}-L_{2,k}\Delta{\mathbf{\dot{X}}}_{k+1}-\beta\mathcal{C}^{\#}\mathcal{B}\Delta\mathbf{\dot{W}}_{k+1},\quad\mathbf{\dot{D}}_{\mathbf{\dot{\Lambda}}_{k+1}}:=\frac{1}{\beta\mu}\Delta\mathbf{\dot{\Lambda}}_{k+1},\,$
$\mathbf{\dot{D}}_{\mathbf{\dot{W}}_{k+1}}$ $:=\nabla g(\mathbf{\dot{W}}_{k+1})-\nabla g(\mathbf{\dot{W}}_{k})+\mathcal{B}^{\#}\Delta\mathbf{\dot{\Lambda}}_{k+1}-L_{1,k}\Delta\mathbf{\dot{W}}_{k+1}.$
Furthermore, there exists $\rho>0$ such that
@@
$-\frac{1}{\tau^{2}}\mathcal{A}^{\#}\mathcal{P}^{\#}_{\Omega}(\mathcal{P}_{\Omega}(\mathcal{A}{\mathbf{\dot{X}}}_{k+1}-\mathbf{\dot{Y}}))-\beta\mathcal{C}^{\#}(\mathcal{C}({\mathbf{\dot{X}}}_{k+1})+\mathcal{B}(\mathbf{\dot{W}}_{k})-\mathbf{\dot{B}}+\mathbf{\dot{\Lambda}}_{k}/\beta)-L_{1,k}\Delta{\mathbf{\dot{X}}}_{k+1}\in\partial f({\mathbf{\dot{X}}}_{k+1}),$
@@
$\mathbf{\dot{D}}_{{\mathbf{\dot{X}}}_{k+1}}=-\mathcal{C}^{\#}\Delta\mathbf{\dot{\Lambda}}_{k+1}-L_{1,k}\Delta{\mathbf{\dot{X}}}_{k+1}+\beta\mathcal{C}^{\#}\mathcal{B}\Delta\mathbf{\dot{W}}_{k+1}\in\partial_{{\mathbf{\dot{X}}}}\mathcal{L}_{\beta}({\mathbf{\dot{X}}}_{k+1},\mathbf{\dot{W}}_{k+1},\mathbf{\dot{\Lambda}}_{k+1}).$ (39)

### v2

@@ 4.1 Global Convergence for Algorithm 1
Suppose A2,A3 in Assumption 4.1 hold, the sequence $\{({\mathbf{\dot{X}}}_{k},\mathbf{\dot{W}}_{k},\mathbf{\dot{\Lambda}}_{k})\}_{k\geq 0}$ generated by (59) satisfies:
$\frac{1}{\beta\mu}\|\Delta\mathbf{\dot{\Lambda}}_{k+1}\|^{2}_{\mathrm{F}}\leq\varsigma_{0}\|\Delta\mathbf{\dot{W}}_{k}\|^{2}_{\mathrm{F}}+\varsigma_{1}\|\Delta\mathbf{\dot{W}}_{k+1}\|^{2}_{\mathrm{F}}+\varsigma_{2}\|\mathcal{B}^{\#}(\Delta\mathbf{\dot{\Lambda}}_{k})\|^{2}_{\mathrm{F}}-\varsigma_{2}\|\mathcal{B}^{\#}(\Delta\mathbf{\dot{\Lambda}}_{k+1})\|^{2}_{\mathrm{F}},$ (30)
$\mathcal{L}_{\beta}({\mathbf{\dot{X}}}_{k+1},\mathbf{\dot{W}}_{k+1},\mathbf{\dot{\Lambda}}_{k+1})+\|\Delta\mathbf{\dot{W}}_{k+1}\|_{B_{k}-r\varsigma_{1}\mathbf{I}}^{2}+\frac{L_{k,1}}{2}\|\Delta{\mathbf{\dot{X}}}_{k+1}\|^{2}_{\mathrm{F}}+\frac{r-1}{\beta\mu}\|\Delta\mathbf{\dot{\Lambda}}_{k+1}\|^{2}_{\mathrm{F}}$
$+r\varsigma_{2}\|\mathcal{B}^{\#}\Delta\mathbf{\dot{\Lambda}}_{k+1}\|^{2}_{\mathrm{F}}\leq\mathcal{L}_{\beta}({\mathbf{\dot{X}}}_{k},\mathbf{\dot{W}}_{k},\mathbf{\dot{\Lambda}}_{k})+r\varsigma_{0}\|\Delta\mathbf{\dot{W}}_{k}\|^{2}_{\mathrm{F}}+r\varsigma_{2}\|\mathcal{B}^{\#}(\Delta\mathbf{\dot{\Lambda}}_{k})\|^{2}_{\mathrm{F}},$ (31)
where $r>1,\varsigma_{0},\varsigma_{1}$ are defined in (27), and $\varsigma_{2}:=\frac{|1-\mu|}{\beta\mu\lambda_{+}^{\mathcal{B}^{\#}\mathcal{B}}\rho(\mu)}.$
@@
$\mathbf{\dot{G}}_{k+1}:=-L_{k,2}\Delta\mathbf{\dot{W}}_{k+1}-\nabla g(\mathbf{\dot{W}}_{k}).$ (32)
It follows that $\Delta\mathbf{\dot{G}}_{k+1}=L_{2,k-1}\Delta\mathbf{\dot{W}}_{k}-L_{k,2}\Delta\mathbf{\dot{W}}_{k+1}+\nabla g(\mathbf{\dot{W}}_{k-1})-\nabla g(\mathbf{\dot{W}}_{k}).$ By the triangle inequality, the following inequality holds:
$\|\Delta\mathbf{\dot{G}}_{k+1}\|_{\mathrm{F}}\leq\|\nabla g(\mathbf{\dot{W}}_{k-1})-\nabla g(\mathbf{\dot{W}}_{k})\|_{\mathrm{F}}+L_{k,2}\|\Delta\mathbf{\dot{W}}_{k+1}\|_{\mathrm{F}}+L_{2,k-1}\|\Delta\mathbf{\dot{W}}_{k}\|_{\mathrm{F}}.$
By A2 in Assumption 4.1, $\nabla g(\mathbf{\dot{W}})$ is $L_{g}$ Lipschitz continuous and $q_{2}=\sup_{k\geq 0}L_{k,2}<\infty,$ we have $\|\Delta\mathbf{\dot{G}}_{k+1}\|_{\mathrm{F}}\leq(L_{g}+q_{2})\|\Delta\mathbf{\dot{W}}_{k}\|_{\mathrm{F}}+q_{2}\|\Delta\mathbf{\dot{W}}_{k+1}\|_{\mathrm{F}}.$ Hence it follows that:
@@
Let $\mathcal{R}({\mathbf{\dot{X}}},\mathbf{\dot{W}},\mathbf{\dot{\Lambda}},\mathbf{\dot{W}}^{\prime},\mathbf{\dot{\Lambda}}^{\prime}):=\mathcal{L}_{\beta}({\mathbf{\dot{X}}},\mathbf{\dot{W}},\mathbf{\dot{\Lambda}})+r\varsigma_{0}\|\mathbf{\dot{W}}-\mathbf{\dot{W}}^{\prime}\|^{2}_{\mathrm{F}}+r\varsigma_{2}\|\mathcal{B}^{\#}(\mathbf{\dot{\Lambda}}-\mathbf{\dot{\Lambda}}^{\prime})\|^{2}_{\mathrm{F}}$. The merit function $\mathcal{R}_{k}$ is defined by
$\mathcal{R}_{k}:=\mathcal{R}({\mathbf{\dot{X}}}_{k},\mathbf{\dot{W}}_{k},\mathbf{\dot{\Lambda}}_{k},\mathbf{\dot{W}}_{k-1},\mathbf{\dot{\Lambda}}_{k-1})=\mathcal{L}_{\beta}({\mathbf{\dot{X}}}_{k},\mathbf{\dot{W}}_{k},\mathbf{\dot{\Lambda}}_{k})+r\varsigma_{0}\|\Delta\mathbf{\dot{W}}_{k}\|^{2}_{\mathrm{F}}+r\varsigma_{2}\|\mathcal{B}^{\#}\Delta\mathbf{\dot{\Lambda}}_{k}\|^{2}_{\mathrm{F}}.$ (36)
@@
$\leq\mathcal{R}_{k+1}+\|\Delta\mathbf{\dot{W}}_{k+1}\|_{B_{k}-r(\varsigma_{1}+\varsigma_{0})\mathbf{I}}^{2}+\frac{L_{k,1}}{2}\|\Delta{\mathbf{\dot{X}}}_{k+1}\|^{2}_{\mathrm{F}}+\frac{r-1}{\beta\mu}\|\Delta\mathbf{\dot{\Lambda}}_{k+1}\|^{2}_{\mathrm{F}}\leq\mathcal{R}_{k}\leq\mathcal{R}_{k_{0}},$
@@
$-\frac{1}{2\beta}\left\|\mathbf{\dot{\Lambda}}_{k+1}\right\|_{\mathrm{F}}^{2}+(r\varsigma_{0}+a)\|\Delta\mathbf{\dot{W}}_{k+1}\|_{\mathrm{F}}^{2}+a(\|\Delta{\mathbf{\dot{X}}}_{k+1}\|_{\mathrm{F}}^{2}+\|\Delta\mathbf{\dot{\Lambda}}_{k+1}\|_{\mathrm{F}}^{2})+r\varsigma_{2}\|\mathcal{B}^{\#}\Delta\mathbf{\dot{\Lambda}}_{k+1}\|_{\mathrm{F}}^{2}\leq\mathcal{R}_{k_{0}}.$
@@
It follows from the convexity of $\|\cdot\|^{2}$ that $\lambda_{+}^{\mathcal{B}^{\#}\mathcal{B}}\mu^{2}\|\mathbf{\dot{\Lambda}}_{k+1}\|_{\mathrm{F}}^{2}\leq\frac{\mu^{2}}{\rho(\mu)}\|\mathbf{\dot{G}}_{k+1}\|_{\mathrm{F}}^{2}+|1-\mu|\|\mathcal{B}^{\#}\Delta\mathbf{\dot{\Lambda}}_{k+1}\|_{\mathrm{F}}^{2}.$ According to definition (32), we have $\|\mathbf{\dot{G}}_{k+1}\|_{\mathrm{F}}^{2}\leq 2q_{2}^{2}\|\Delta\mathbf{\dot{W}}_{k+1}\|_{\mathrm{F}}^{2}+2\|\nabla g(\mathbf{\dot{W}}_{k})\|_{\mathrm{F}}^{2}\leq 2(q_{2}+L_{g})^{2}\|\Delta\mathbf{\dot{W}}_{k+1}\|_{\mathrm{F}}^{2}+2\|\nabla g(\mathbf{\dot{W}}_{k+1})\|_{\mathrm{F}}^{2}.$ Hence it follows that:
$-\frac{1}{2\beta}\|\mathbf{\dot{\Lambda}}_{k+1}\|_{\mathrm{F}}^{2}\geq-\varsigma_{3}\|\nabla g(\mathbf{\dot{W}}_{k+1})\|_{\mathrm{F}}^{2}-\varsigma_{4}\|\Delta\mathbf{\dot{W}}_{k+1}\|_{\mathrm{F}}^{2}-\varsigma_{5}\|\mathcal{B}^{\#}\Delta\mathbf{\dot{\Lambda}}_{k+1}\|_{\mathrm{F}}^{2},$ (40)
@@
$\varsigma_{3}:=\frac{1}{\beta\rho(\mu)\lambda_{+}^{\mathcal{B}^{\#}\mathcal{B}}},\quad\varsigma_{4}:=\frac{(q_{2}+L_{g})^{2}}{\beta\rho(\mu)\lambda_{+}^{\mathcal{B}^{\#}\mathcal{B}}},\quad\varsigma_{5}:=\frac{|1-\mu|}{2\beta\mu^{2}\lambda_{+}^{\mathcal{B}^{\#}\mathcal{B}}}.$
@@
$f({\mathbf{\dot{X}}}_{k+1})+(1-\kappa)g(\mathbf{\dot{W}}_{k+1})+\frac{1}{2\tau^{2}}\|\mathcal{P}_{\Omega}(\mathcal{A}({\mathbf{\dot{X}}}_{k}))-\mathcal{P}_{\Omega}(\mathbf{\dot{Y}})\|_{\mathrm{F}}^{2}+\frac{\beta}{2}\left\|\mathcal{C}({\mathbf{\dot{X}}}_{k+1})+\mathcal{B}(\mathbf{\dot{W}}_{k+1})-\mathbf{\dot{B}}+\frac{\mathbf{\dot{\Lambda}}_{k+1}}{\beta}\right\|_{\mathrm{F}}^{2}$ (41)
$+a(\|\Delta{\mathbf{\dot{X}}}_{k+1}\|_{\mathrm{F}}^{2}+\|\Delta\mathbf{\dot{\Lambda}}_{k+1}\|_{\mathrm{F}}^{2})+(r\varsigma_{0}+a-\varsigma_{4})\|\Delta\mathbf{\dot{W}}_{k+1}\|_{\mathrm{F}}^{2}+(r\varsigma_{2}-\varsigma_{5})\|\mathcal{B}^{\#}\Delta\mathbf{\dot{\Lambda}}_{k+1}\|_{\mathrm{F}}^{2}$
$\leq\mathcal{R}_{k_{0}}-\inf_{\mathbf{\dot{W}}}\left\{\kappa g(\mathbf{\dot{W}})-\varsigma_{3}\|\nabla g(\mathbf{\dot{W}})\|_{\mathrm{F}}^{2}\right\},$
where $\kappa\in(0,1).$ According to (25) in Assumption A2, setting $\mathbf{\dot{W}}_{1}=\mathbf{\dot{W}}-\delta\nabla g(\mathbf{\dot{W}})$ and $\mathbf{\dot{W}}_{2}=\mathbf{\dot{W}}$, it follows that $\kappa g(\mathbf{\dot{W}}_{k}-\delta\nabla g(\mathbf{\dot{W}}_{k}))\leq\kappa g(\mathbf{\dot{W}}_{k})-\kappa(\delta-\frac{L_{g}\delta^{2}}{2})\|\nabla g(\mathbf{\dot{W}}_{k})\|_{\mathrm{F}}^{2}.$ Since $g$ is bounded from below, there exist $M$ such that
$-M<\inf\{\kappa g(\mathbf{\dot{W}})-\kappa(\delta-\frac{L_{g}\delta^{2}}{2})\|\nabla g(\mathbf{\dot{W}})\|_{\mathrm{F}}^{2}:\mathbf{\dot{W}}\in\mathbb{H}^{m\times n}\}.$ (42)
We choose $\delta=\frac{1}{L_{g}}$. According to Assumption A5, we have $\varsigma_{3}<\frac{\kappa}{2L_{g}}=\kappa(\delta-\frac{L_{g}\delta^{2}}{2}).$ Since $r>1$ and $\mu\in(0,2),$ according to the denifition of $\varsigma_{0},\varsigma_{2},\varsigma_{4}$ and $\varsigma_{5}$, it holds that $r\varsigma_{2}-\varsigma_{5}>0$ and $r\varsigma_{0}+a-\varsigma_{4}>0$. It follows from (41) that
$f({\mathbf{\dot{X}}}_{k+1})+(1-\kappa)g(\mathbf{\dot{W}}_{k+1})+\frac{1}{2\tau^{2}}\|\mathcal{P}_{\Omega}(\mathcal{A}({\mathbf{\dot{X}}}_{k}))-\mathcal{P}_{\Omega}(\mathbf{\dot{Y}})\|_{\mathrm{F}}^{2}$ (43)
@@
Since $f$ is coercive, the sequence $\{{\mathbf{\dot{X}}}_{k}\}_{k\geq k_{0}}$ is bounded and hence $\{\mathcal{C}{\mathbf{\dot{X}}}_{k}\}_{k\geq k_{0}}$ is bounded. According to (43), $\Delta\mathbf{\dot{\Lambda}}_{k+1}$ is bounded. By the $\mathbf{\dot{\Lambda}}$ update in (59), we have $\mathcal{B}\mathbf{\dot{W}}_{k+1}=\frac{1}{\beta\mu}\Delta\mathbf{\dot{\Lambda}}_{k+1}-\mathcal{C}{\mathbf{\dot{X}}}_{k+1}+\mathbf{\dot{B}}.$ Since $\{\mathcal{C}{\mathbf{\dot{X}}}_{k}\}_{k\geq k_{0}}$ is bounded, if $\mathcal{B}^{\#}\mathcal{B}$ is full rank or $g$ is coercive, $\{\mathbf{\dot{W}}_{k}\}_{k\geq k_{0}}$ is bounded. From the fact that $\frac{\beta}{2}\left\|\mathcal{C}({\mathbf{\dot{X}}}_{k+1})+\mathcal{B}(\mathbf{\dot{W}}_{k+1})-\mathbf{\dot{B}}+\frac{\mathbf{\dot{\Lambda}}_{k+1}}{\beta}\right\|_{\mathrm{F}}^{2}$ is bounded, it follows that $\{\mathbf{\dot{\Lambda}}_{k}\}_{k\geq k_{0}}$ is bounded. As a consequence, $\{{\mathbf{\dot{X}}}_{k}\}_{k\geq 1},\{\mathbf{\dot{W}}_{k}\}_{k\geq 1},\{\mathbf{\dot{\Lambda}}_{k}\}_{k\geq 1}$ is bounded. ∎
@@
$\mathbf{\dot{D}}_{{\mathbf{\dot{X}}}_{k+1}}$ $:=\mathcal{C}^{\#}\Delta\mathbf{\dot{\Lambda}}_{k+1}-L_{k,1}\Delta{\mathbf{\dot{X}}}_{k+1}-\beta\mathcal{C}^{\#}\mathcal{B}\Delta\mathbf{\dot{W}}_{k+1},\quad\mathbf{\dot{D}}_{\mathbf{\dot{\Lambda}}_{k+1}}:=\frac{1}{\beta\mu}\Delta\mathbf{\dot{\Lambda}}_{k+1},\,$
$\mathbf{\dot{D}}_{\mathbf{\dot{W}}_{k+1}}$ $:=\nabla g(\mathbf{\dot{W}}_{k+1})-\nabla g(\mathbf{\dot{W}}_{k})+\mathcal{B}^{\#}\Delta\mathbf{\dot{\Lambda}}_{k+1}-L_{k,2}\Delta\mathbf{\dot{W}}_{k+1}.$
Furthermore, there exists $\pi>0$ such that
@@
$-\frac{1}{\tau^{2}}\mathcal{A}^{\#}\mathcal{P}^{\#}_{\Omega}(\mathcal{P}_{\Omega}(\mathcal{A}{\mathbf{\dot{X}}}_{k+1}-\mathbf{\dot{Y}}))-\beta\mathcal{C}^{\#}(\mathcal{C}({\mathbf{\dot{X}}}_{k+1})+\mathcal{B}(\mathbf{\dot{W}}_{k})-\mathbf{\dot{B}}+\mathbf{\dot{\Lambda}}_{k}/\beta)-L_{k,1}\Delta{\mathbf{\dot{X}}}_{k+1}\in\partial f({\mathbf{\dot{X}}}_{k+1}),$
@@
$\mathbf{\dot{D}}_{{\mathbf{\dot{X}}}_{k+1}}=-\mathcal{C}^{\#}\Delta\mathbf{\dot{\Lambda}}_{k+1}-L_{k,1}\Delta{\mathbf{\dot{X}}}_{k+1}+\beta\mathcal{C}^{\#}\mathcal{B}\Delta\mathbf{\dot{W}}_{k+1}\in\partial_{{\mathbf{\dot{X}}}}\mathcal{L}_{\beta}({\mathbf{\dot{X}}}_{k+1},\mathbf{\dot{W}}_{k+1},\mathbf{\dot{\Lambda}}_{k+1}).$ (45)